### v1

# Dynamic compensation for pump-induced frequency shift in Kerr-cat qubit initialization

Abstract
The noise-biased Kerr-cat qubit is an attractive candidate for fault-tolerant quantum computation; however, its initialization faces challenges due to the squeezing pump-induced frequency shift (PIFS). Here, we propose and demonstrate a dynamic compensation method to mitigate the effect of PIFS during the Kerr-cat qubit initialization. Utilizing a novel nonlinearity-engineered triple-loop SQUID device, we realize a stabilized Kerr-cat qubit and validate the advantages of the dynamic compensation method by improving the initialization fidelity from 57% to 78%, with a projected fidelity of 91% after excluding state preparation and measurement errors. Our results not only advance the practical implementation of Kerr-cat qubits, but also provide valuable insights into the fundamental adiabatic dynamics of these systems. This work paves the way for scalable quantum processors that leverage the bias-preserving properties of Kerr-cat qubits.
Introduction.- Bias-preserving cat qubits Mirrahimi et al. (2014); Leghtas et al. (2015), which encode quantum information in Schrödinger cat states, have emerged as a promising platform for fault-tolerant quantum computation Aliferis and Preskill (2008); Aliferis et al. (2009); Webster et al. (2015); Chamberland et al. (2022). These qubits offer exponential suppression of bit-flip errors Mirrahimi et al. (2014); Lescanne et al. (2020); Puri et al. (2017) and reduced resource overhead for quantum error correction. The bias-preserving property of cat qubits is particularly advantageous for implementing surface codes Fowler et al. (2012); Bonilla Ataides et al. (2021); Tuckett et al. (2018, 2019, 2020); Darmawan et al. (2021), repetition codes Shor (1995); Steane (1996); Guillaud and Mirrahimi (2019, 2021); Gouzien et al. (2023), and error syndrome detection Puri et al. (2019); Ding et al. (2024). Exciting experimental progress has been achieved recently, with cat qubit lifetimes exceeding 100 seconds without quantum control Berdou et al. (2023) and 10 seconds with quantum control Réglade et al. (2024).
Two main approaches exist for realizing cat qubits: dissipative cat qubits Mirrahimi et al. (2014); Lescanne et al. (2020) and Kerr-cat qubits Puri et al. (2017); Grimm et al. (2020). Kerr-cat qubits are particularly appealing due to their potential for high-speed gate operations Xu et al. (2022). However, their initialization, which typically involves adiabatically ramping up a two-photon pump to map the Fock states to the cat state manifold Puri et al. (2017); Grimm et al. (2020), faces a significant challenge: the pump-induced frequency shift (PIFS). This shift, arising from either charge Grimm et al. (2020) or flux Lu et al. (2023); Iyama et al. (2024) pumping, can cause decoherence and state leakage during the initialization Grimm et al. (2020); Puri et al. (2020). Previous attempts to compensate for PIFS rely on applying a static detuning between the pump and qubit frequencies Grimm et al. (2020). While this approach works for small PIFS, it fails to account for the dynamic nature of PIFS during the ramping-up process, leading to imperfect initialization. Moreover, as the pursuit of better bias-preserving properties necessitates larger cat amplitudes, the effect of PIFS becomes increasingly pronounced and demands a more sophisticated compensation strategy.
In this work, we introduce a dynamic compensation method that continuously adjusts the pump frequency to match the instantaneous PIFS during the adiabatic initialization. We experimentally demonstrate the advantage of this dynamic compensation for a Kerr-cat qubit using a novel Nonlinearity-Engineered Triple-loop SQUID (NETS) device Hua et al. . This device not only holds promise for realizing bias-preserving CNOT gates Puri et al. (2020); Chamberland et al. (2022), but also offers independent tuning of nonlinearities compared with other Kerr-cat qubit candidates such as SNAIL Frattini et al. (2017, 2018); Sivak et al. (2019), ATS Lescanne et al. (2020), Josephson ring modulator Bergeal et al. (2010); Roch et al. (2012), and a series of asymmetric DC SQUIDs Iyama et al. (2024). Our theoretical and experimental results showcase the superiority of the dynamic compensation approach, particularly for large PIFS values. Our demonstration paves the way for high-fidelity Kerr-cat qubit operations and advances the field towards fault-tolerant quantum computation.
Figure 1: (a) Bloch sphere representation of the Kerr-cat qubit encoding space. Here, |±$\left|\pm\alpha\right\rangle$ denotes the coherent states and $\left|C_{\alpha}^{k}\right\rangle=(\left|\alpha\right\rangle+k\left|-\alpha\right\rangle)/\sqrt{2}$ for $k=+,-,+\mathrm{i},-\mathrm{i}$. (b) Energy levels of the system as a function of $\Delta(0)$, the mode detuning at $t=0$ in Eq. 1 with no external pump. (c) Solid lines: evolution of the energy levels for an ideal case with $\Delta\equiv 0$ when ramping up the pump amplitude $\epsilon_{2}$ to a final value of $\epsilon_{\mathrm{2,KC}}$. Dashed lines: the removal of the degeneracy due to the pump-induced frequency shift (PIFS). (d) Evolution of the energy levels for a static compensation with a fixed pump detuning $\Delta_{\mathrm{as}}(t)\equiv 1.5K$. At $\epsilon_{2}=0$, $\Delta(0)=1.5K$, corresponding to the right dashed line in (b). In (c) and (d), the purple arrows indicate adiabatic evolution paths that map the vacuum state to $\left|C_{\alpha}^{+}\right\rangle$ and to $\left|\psi_{\mathrm{e,1}}^{+}\right\rangle\approx[\hat{D}(\alpha)+\hat{D}(-\alpha)]\left|1\right\rangle$, respectively.
Principle.- The Kerr-cat qubit can be realized with a single-bosonic mode Hamiltonian ($\hbar=1$Puri et al. (2017); Grimm et al. (2020)
$\hat{H}_{\mathrm{cat}}=\Delta(t)a^{\dagger}a-K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\varepsilon_{2}(t)\hat{a}^{\dagger 2}+\varepsilon_{2}^{*}(t)\hat{a}^{2},$ (1)
where $\hat{a}^{\dagger}(\hat{a})$ is the creation (annihilation) operator of the mode, $\Delta$ is the relative frequency detuning of the mode, $K$ is the Kerr nonlinearity, and $\varepsilon_{2}$ is the pump amplitude. In an ideal experimental configuration with $\Delta=0$, the bosonic mode has a stabilized two-dimensional Hilbert space spanned by |±$\left|\pm\alpha\right\rangle$, i.e., the coherent states with an average photon number $|\alpha|^{2}=\varepsilon_{2}/K$. Consequently, the superposition states $\left|C_{\alpha}^{k}\right\rangle=(\left|\alpha\right\rangle+k\left|-\alpha\right\rangle)/\sqrt{2}$ for $k=+,-,+\mathrm{i},-\mathrm{i}$ are also stabilized states in the mode. Figure 1(a) presents the Bloch sphere of the Kerr-cat qubit by treating $\left|C_{\alpha}^{\pm}\right\rangle$ as the computational basis. The Kerr-cat qubit has the advantage of being bias-preserving, with the states along the X-axis having an enhanced lifetime that increases exponentially with the Kerr-cat size $|\alpha|^{2}$ Puri et al. (2017, 2020).
Figure 2: (a) Circuit schematic of the device. The Nonlinearity-Engineered Triple-loop SQUID (NETS) is shunted with a large capacitor, similar to that of a transmon. (b) Circuit schematic of the NETS. $I_{2}$ is the flux pumping port. The loop areas are meticulously designed such that the flux ratios are 1:-1:3, and the ratio of the small (pink) to the large (blue) Josephson junctions is $r\approx 0.16$. More details and parameters about the device are provided in Ref. SM . (c) Photograph of the NETS. (d) Illustration comparing the dynamic and static compensation methods. The envelope $\varepsilon_{2}(t)$ follows a tanh function. In the static method, the pump frequency remains fixed over time, whereas in the dynamic method, the pump frequency varies with time $\Delta_{\mathrm{as}}(t)\propto\varepsilon^{2}_{2}(t)$. (e-l) Comparison of the ideal and experimental Wigner functions for two typical PIFS. (e-h) The first two even-parity eigenstates for $|\alpha|^{2}\approx 1.1$ and 1.9, with $\Delta_{\mathrm{as,KC}}=3.4\,\mathrm{MHz}<K$ and $\Delta_{\mathrm{as,KC}}=9.7\,\mathrm{MHz}>K$. (i-l) The experimental results based on the dynamic and static methods. The color contrast is mainly limited by the decoherence of the transmon.
The initialization or encoding of the Kerr-cat qubit can be realized by ramping up the amplitude of the squeezing pump $\varepsilon_{2}$. As shown by the solid lines in Fig. 1(c), the eigenstates of the Hamiltonian evolve from the Fock state basis $\left|0\right\rangle$, $\left|1\right\rangle$} to the computational basis of the Kerr-cat qubit following the traces as $\varepsilon_{2}$ increases. However, in practical experiments there are parasitic effects on the mode when applying pumps to it. Specifically, we encounter a dynamic PIFS $\Delta_{\mathrm{PI}}(t)=\gamma\varepsilon_{2}^{2}(t)$ with a coefficient $\gamma$. Consequently, the corresponding detuning varies with the pump as $\Delta=\Delta_{\mathrm{as}}-\Delta_{\mathrm{PI}}$, where $\Delta_{\mathrm{as}}=\omega_{\mathrm{a}}-\omega_{\mathrm{s}}/2$ is the detuning between the external squeezing pump ($\omega_{\mathrm{s}}$) and the bare mode ($\omega_{\mathrm{a}}$). We define the pump amplitude at the end of the ramping-up time $t_{\mathrm{up}}$ as $|\varepsilon_{2}(t_{\mathrm{up}})|=\varepsilon_{\mathrm{2,KC}}$ and the detuning $\Delta_{\mathrm{as}}(t_{\mathrm{up}})=\Delta_{\mathrm{as,KC}}$. Then, $\Delta_{\mathrm{PI}}(0)=0$ and $\Delta_{\mathrm{PI}}(t_{\mathrm{up}})=\Delta_{\mathrm{PI,KC}}=\gamma\varepsilon^{2}_{\mathrm{2,KC}}$. Here the subscript KC denotes the parameters for preserving the Kerr-cat qubit.
If the PIFS is not compensated, i.e., $\Delta_{\mathrm{as}}\equiv 0$, then $\Delta=0$ holds true only at $t=0$. As $\Delta$ increases when ramping up the pump, the corresponding eigenenergies of the Hamiltonian become non-degenerate, as shown by the dashed lines in Fig. 1(c) with $\Delta_{\mathrm{PI,KC}}=0.8K$ as an example. Additionally, the prepared Kerr-cat qubit at $t>t_{\mathrm{up}}$ experiences a non-zero $\Delta$, which induces a Z-rotation of the Kerr-cat qubit [orange arrow in Fig. 1(a)] with a precession rate $\Omega_{Z}=4\Delta|\alpha|^{2}\exp(-2|\alpha|^{2})$  Grimm et al. (2020); Puri et al. (2020). This rotation causes the state in the well-protected axis to pass through the not-protected axis, violating the principle of bias-preserving qubits and significantly shortening the coherence time.
A simple solution to this problem is to introduce a static detuning of the pump to compensate $\Delta_{\mathrm{PI,KC}}$ after the initialization, i.e., $\Delta_{\mathrm{as}}(t)\equiv\Delta_{\mathrm{as,KC}}=\Delta_{\mathrm{PI,KC}}$. However, this static compensation lifts the degeneracy of Fock states at the beginning of the initialization due to $\Delta(0)=\Delta_{\mathrm{as,KC}}$. As illustrated in Fig. 1(b), the eigenenergies of the eigenstates (Fock states) at $t=0$ vary with $\Delta(0)$, resulting in a rearrangement of the order of the Fock states. Similar results are also discussed in Ref. Yanagimoto et al. (2019). When ramping up $\varepsilon_{2}$, as shown in Fig. 1(d) for $\Delta_{\mathrm{PI,KC}}=1.5K$ as an example, the evolution of eigenstates is significantly different from that in Fig. 1(c).
Two potential issues arise for such a static compensation. First, the traces of the eigenstates change significantly, and we should choose higher photon number states for adiabatic initialization of the target cat states. Otherwise, we end up with a leakage state [purple arrow in Fig. 1(d)]. Second, the energy gaps between different traces are reduced as compared to the ideal cases in Fig. 1(c), imposing stringent adiabatic ramping rate, i.e., $t_{\mathrm{up}}\gg 1/\delta_{E}$, where $\delta_{E}$ is the smallest energy gap Dirac and Rutherford (1925). Considering the relationship $\Delta_{\mathrm{PI,KC}}=\gamma|\alpha|^{4}K^{2}$, such effects cannot be neglected when attempting to achieve excellent bias noise properties by increasing $|\alpha|^{2}$.
Therefore, we propose an alternative approach that dynamically varies the pump frequency to compensate the PIFS while maintaining $\Delta=0$ throughout the entire ramping process. This approach effectively recovers the ideal eigenenergy traces shown in Fig. 1(c), enabling high-fidelity initialization of the Kerr-cat qubit.
Experiments.- The Kerr-cat qubit is implemented using a novel NETS device with its schematic circuit shown in Fig. 2(a). An ancillary transmon qubit is introduced for characterizing the Kerr-cat qubit. The NETS consists of three loops [Fig. 2(b)], holding the advantage of independently tuning the nonlinearities of different orders by controlling the external flux through three current lines $I_{1}$, $I_{2}$, and $I_{3}$ Hua et al. ; SM . Figure 2(c) depicts the fabricated NETS device, with the ratios of the junctions and loop areas being meticulously designed. By biasing the NETS at $(\phi_{\mathrm{e1}},\phi_{\mathrm{e2}},\phi_{\mathrm{e3}})=(-0.08,1.08,0)\pi$ and applying a flux pump via $I_{2}$ line, the Hamiltonian of Eq. 1 for the Kerr-cat qubit can be realized SM .
Without an external pump, the NETS functions as a nonlinear oscillator at a frequency of $\omega_{\mathrm{a}}/2\pi=5.6\,\mathrm{GHz}$ with a Kerr nonlinearity $K/2\pi=6.9\,\mathrm{MHz}$. The amplitude damping and transverse coherence times are $T_{\mathrm{1N}}=6.0\,\mathrm{\mu s}$ and $T_{\mathrm{2N}}=3.0\,\mathrm{\mu s}$, respectively. The Kerr-cat qubit is realized by applying a flux pump at $\omega_{\mathrm{s}}/2\pi\approx 11.2\,\mathrm{GHz}$ with an amplitude of $\varepsilon_{\mathrm{2,KC}}/2\pi=9.7\,\mathrm{MHz}$ to the NETS. By initializing the NETS in the state $(\left|0\right\rangle+\left|1\right\rangle)/\sqrt{2}$, the Kerr-cat qubit can be prepared in the $\left|\alpha\right\rangle$ state on the equator of the Bloch sphere. The precession due to PIFS is measured by varying the holding time of NETS in the Kerr-cat states, allowing the calibration of $\Delta_{\mathrm{as,KC}}$ SM . Furthermore, the Rabi drive of the Kerr-cat qubit is realized by a charge drive on the NETS, with a Rabi frequency along the X-axis given by $\Omega_{X}=\mathrm{Re}(4\epsilon_{\mathrm{x}}\alpha)/\sqrt{1-\exp(-4|\alpha|^{2})}$  Puri et al. (2020). Here, $\epsilon_{\mathrm{x}}$ is the Rabi drive strength. By fitting the Rabi dynamics, the Kerr-cat size is calibrated as $|\alpha|^{2}=1.41$, with a significantly extended damping time of $T_{\mathrm{1,KC}}=13\,\mathrm{\mu s}$ SM .
Next, we systematically investigate the performance of the static and dynamic compensation methods in mapping between the Fock states and Kerr-cat qubit states. Specifically, we focus on the mapping from $\left|0\right\rangle$ to $\left|C_{\alpha}^{+}\right\rangle$, since this mapping is not subject to gate fidelity in the Fock space. Figure 2(d) illustrates the frequency (vertical axis) and amplitude (color) profiles for the ramping-up process, comparing the two methods. Both schemes utilize a tanh function for the amplitude $\varepsilon_{2}$, while the dynamic compensation method employs a tanh-squared function for the frequency, with a ramping-up time $t_{\mathrm{up}}=320~{}\mathrm{ns}$ that satisfies the adiabatic condition $t_{\mathrm{up}}\gg 1/2K$ Grimm et al. (2020). The experimental Wigner functions Sun et al. (2014); Iyama et al. (2024) of the Kerr-cat state with $|\alpha|^{2}\approx 1.1$ and 1.9, using different compensation schemes, are shown in Figs. 2(i)-(l). For a small $|\alpha|^{2}$, where the PIFS is smaller than $K$, both methods yield the same results [Figs. 2(i) and 2(j)] that match the target state $\left|C_{\alpha}^{+}\right\rangle$ [Fig. 2(e)] but are distinct from the potential leakage state $\left|\psi_{e,1}^{+}\right\rangle$ [Fig. 2(f)]. However, when the PIFS is larger than $K$ but smaller than $3K$, i.e., $1.4K$ for $|\alpha|^{2}=1.9$, the dynamic approach outperforms the static approach. As shown in Fig. 2(k), the dynamic compensation method yields a state that more closely resembles the ideal eigenstate [Fig. 2(g)]. In contrast, the static compensation method leads the state [Fig. 2(l)] to deviate significantly from the ideal case and appear more similar to the leakage state $\left|\psi_{\mathrm{e,1}}^{+}\right\rangle$ [Fig. 2(h)]. This confirms that the static compensation fails to account for the time-dependent nature of the PIFS, leading to imperfect initialization and increased leakage as the PIFS grows larger.
Figure 3: (a) Experimental process fidelities of the combined ramping-up and ramping-down operations for different pump strengths, corresponding to various compensation frequencies ($\Delta_{\mathrm{as,KC}}$) and Kerr-cat sizes ($|\alpha|^{2}$), using the dynamic and static compensation methods. Vertical dashed lines: the coalescence of Fock state energies around $\Delta_{\mathrm{as,KC}}=K$ and $3K$. The data points are averaged over five repeated measurements, with error bars indicating the standard deviation. (b) and (c) Evolution of the energy levels for the static compensation method with $\Delta_{\mathrm{as,KC}}=K$ and $3K$, respectively, illustrating the Landau-Zener transition for Fock states $\{\left|0\right\rangle,\left|2\right\rangle\}$ and $\{\left|1\right\rangle,\left|3\right\rangle\}$.
The fidelities of the mapping based on the two schemes are further characterized by implementing a symmetric ramping-down process to remap the Kerr-cat states back to the Fock space. This whole process is ideally an identity channel, and can be directly characterized by performing quantum process tomography (QPT) Nielsen and Chuang (2010) in the Fock state basis when the pump is off. Figure 3(a) presents the process fidelity for different pump amplitudes, i.e., different PIFS and $|\alpha|^{2}$, obtained using the two schemes. It demonstrates that the dynamic compensation method consistently outperforms the static compensation method across the entire range of PIFS values. In particular, when $\Delta_{\mathrm{as,KC}}$ is near $K$ or $3K$, the static compensation method shows strong oscillations of the fidelity. This can be explained by the collapse of the energy levels of the Fock states with respect to the compensated detunings, and the condition for the adiabatic evolution is violated. This leads to the observation of Landau-Zener transitions Landau (1932); Zener and Fowler (1932): an initial Fock state can be mapped to two stabilized states of the pumped NETS when the energy gap between the Fock state and other higher levels becomes small. Due to the energy difference of the stabilized states, the accumulated phase leads to an interference between the two states when mapping the states back to the Fock state basis, exhibiting the sharp oscillations. The evolutions of the energy levels corresponding to these specific cases are shown in Figs. 3(b) and 3(c), respectively. It is worth noting that the squeezing operation preserves parity, which explains the absence of transitions around $\Delta_{\mathrm{as,KC}}=2K$. These observations are also validated by numerical simulations based on QuTip Johansson et al. (2012, 2013), showing excellent agreement between the experimental and numerical results.
The static compensation method shows comparable performance to the dynamic compensation approach when $K<\Delta_{\mathrm{as,KC}}<3K$. This can be attributed to the fact that a precisely symmetric ramping-down process can adiabatically map the stabilized NETS state back to the original Fock space, regardless of whether the NETS state is inside or outside the Kerr-cat qubit space. To directly characterize the initialization of Kerr-cat qubits, we perform a direct cat readout Grimm et al. (2020); Ding et al. (2024); SM of the NETS by coherently converting its state to the readout cavity. Measurements along different axes are realized with an X/2 gate implemented through Rabi drives and a Z/2 gate achieved by instantaneously switching off the flux pump Grimm et al. (2020); SM .
Figure 4: (a) Experimental process fidelities of the ramping-up operation for a Kerr-cat qubit, depending on $\Delta_{\mathrm{as,KC}}$ (or $|\alpha|^{2}$), using different compensation methods. Measurements are performed in the Kerr-cat space. The dynamic compensation method performs much better than the static compensation method when the detuning is larger than $K$. (b-d) Four example octahedrons on the Bloch sphere. The corresponding locations in (a) are indicated by arrows of different colors, along with solid and hollow arrowheads.
Figure 4(a) shows the experimental QPT results (solid markers) for different $\Delta_{\mathrm{as,KC}}$, confirming several key insights into the performance of the dynamic and static methods. When $\Delta_{\mathrm{as,KC}}/K<1$, both methods exhibit comparable initialization fidelities due to the negligible leakage outside the Kerr-cat qubit space. However, as the PIFS increases beyond $K$, the static compensation method experiences a rapid decline in fidelity due to the leakage. In contrast, the dynamic compensation method maintains a high initialization fidelity throughout the entire range of PIFS values. The simulations incorporating the state preparation and measurement (SPAM) errors (solid lines, mainly contributed by the Z/2 gate SM ) closely match the experimental data, confirming our understanding of the system dynamics and the source of imperfections. Therefore, based on a theoretical model, we can project the achievable fidelity by excluding SPAM errors (dashed line), demonstrating the robustness and effectiveness of the dynamic compensation approach. In particular, for $\Delta_{\mathrm{as,KC}}/K=1.4$ and $|\alpha|^{2}=1.9$, the experimental fidelities are 57% and 78% for these two approaches, respectively, corresponding to 52% and 91% when excluding the SPAM errors. Here, the leakage error of the static compensation method is underestimated due to the imperfections in detecting the leakage states Gautier et al. (2022); Ding et al. (2024). We find that the fidelity degrades with $\Delta_{\mathrm{as,KC}}$ due to the increasing decoherence rates with the Kerr-cat size, and better performance is possible by improving the coherence of the NETS. For instance, a Kerr-cat qubit initialization with a fidelity exceeding 99% is anticipated when the NETS has $T_{\mathrm{1N}}=60\,\mathrm{\mu}$s and $T_{\mathrm{2N}}=30\,\mathrm{\mu}$s.
Conclusion.- We have demonstrated high-fidelity initialization of a Kerr-cat qubit using a novel NETS device with a dynamic compensation method. By carefully analyzing eigenstate evolution of the system under adiabatic pump ramping, we clarify the imperfections arising from PIFS and Landau-Zener transitions. Our experimental and theoretical results consistently show that the dynamic compensation approach outperforms the static compensation scheme, significantly suppressing imperfections associated with large PIFS. At a PIFS to Kerr nonlinearity ratio of $1.4$, the measured initialization fidelity improves from 57% to 78%, indicating a projected fidelity of 91% after excluding SPAM errors. Our demonstration with NETS and dynamic compensation paves the way for future advancements in scalable Kerr-cat qubits, enabling their application in quantum information processing and exploration of noise-biased fault-tolerant techniques.
Acknowledgements.
This work was funded by the National Natural Science Foundation of China (Grants No. 11925404, 92165209, 92365301, 12061131011, 92265210, 11890704, 92365206), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300203 and 2021ZD0301800), and the National Key R&D Program (2017YFA0304303). This work was also supported by the Fundamental Research Funds for the Central Universities and USTC Research Funds of the Double First-Class Initiative. This work was partially carried out at the USTC Center for Micro and Nanoscale Research and Fabrication.

### v2

@@
 Here, |±$\left|\pm\alpha\right\rangle$ denotes the coherent states and $\left|C_{\alpha}^{k}\right\rangle=(\left|\alpha\right\rangle+k\left|-\alpha\right\rangle)/\sqrt{2}$ for $k=+,-,+\mathrm{i},-\mathrm{i}$.
-(b) Energy levels of the system as a function of $\Delta(0)$, the mode detuning at $t=0$ in Eq.
+(b) Energy levels of the system as a function of $\Delta(0)$, the detuning at $t=0$ in Eq.
 1 with no external pump.
@@
 Principle.- The Kerr-cat qubit can be realized with a single-bosonic mode Hamiltonian ($\hbar=1$Puri et al. (2017); Grimm et al. (2020)
-$\hat{H}_{\mathrm{cat}}=\Delta(t)a^{\dagger}a-K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\varepsilon_{2}(t)\hat{a}^{\dagger 2}+\varepsilon_{2}^{*}(t)\hat{a}^{2},$ (1)
-where $\hat{a}^{\dagger}(\hat{a})$ is the creation (annihilation) operator of the mode, $\Delta$ is the relative frequency detuning of the mode, $K$ is the Kerr nonlinearity, and $\varepsilon_{2}$ is the pump amplitude.
+$\hat{H}_{\mathrm{cat}}=\Delta(t)a^{\dagger}a-K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\varepsilon_{2}(t)(\hat{a}^{\dagger 2}+\hat{a}^{2}),$ (1)
+where $\hat{a}^{\dagger}(\hat{a})$ is the creation (annihilation) operator of the mode, $\Delta$ is the relative frequency detuning, $K$ is the Kerr nonlinearity, and $\varepsilon_{2}$ is real and represents the pump amplitude.
 In an ideal experimental configuration with $\Delta=0$, the bosonic mode has a stabilized two-dimensional Hilbert space spanned by |±$\left|\pm\alpha\right\rangle$, i.e., the coherent states with an average photon number $|\alpha|^{2}=\varepsilon_{2}/K$.
@@
 Consequently, the corresponding detuning varies with the pump as $\Delta=\Delta_{\mathrm{as}}-\Delta_{\mathrm{PI}}$, where $\Delta_{\mathrm{as}}=\omega_{\mathrm{a}}-\omega_{\mathrm{s}}/2$ is the detuning between the external squeezing pump ($\omega_{\mathrm{s}}$) and the bare mode ($\omega_{\mathrm{a}}$).
-We define the pump amplitude at the end of the ramping-up time $t_{\mathrm{up}}$ as $|\varepsilon_{2}(t_{\mathrm{up}})|=\varepsilon_{\mathrm{2,KC}}$ and the detuning $\Delta_{\mathrm{as}}(t_{\mathrm{up}})=\Delta_{\mathrm{as,KC}}$.
+We define the pump amplitude at the end of the ramping-up time $t_{\mathrm{up}}$ as $\varepsilon_{2}(t_{\mathrm{up}})=\varepsilon_{\mathrm{2,KC}}$ and the detuning $\Delta_{\mathrm{as}}(t_{\mathrm{up}})=\Delta_{\mathrm{as,KC}}$.
 Then, $\Delta_{\mathrm{PI}}(0)=0$ and $\Delta_{\mathrm{PI}}(t_{\mathrm{up}})=\Delta_{\mathrm{PI,KC}}=\gamma\varepsilon^{2}_{\mathrm{2,KC}}$.

### v3

@@
 Here, we propose and demonstrate a dynamic compensation method to mitigate the effect of PIFS during the Kerr-cat qubit initialization.
-Utilizing a novel nonlinearity-engineered triple-loop SQUID device, we realize a stabilized Kerr-cat qubit and validate the advantages of the dynamic compensation method by improving the initialization fidelity from 57% to 78%, with a projected fidelity of 91% after excluding state preparation and measurement errors.
+Utilizing a novel nonlinearity-engineered multi-loop SQUID device, we realize a stabilized Kerr-cat qubit and validate the advantages of the dynamic compensation method by improving the initialization fidelity from 57% to 78%, with a projected fidelity of 91% after excluding state preparation and measurement errors.
 Our results not only advance the practical implementation of Kerr-cat qubits, but also provide valuable insights into the fundamental adiabatic dynamics of these systems.
@@
 (2023); Iyama et al.
+(2024); Hoshi et al.
 (2024) pumping, can cause decoherence and state leakage during the initialization Grimm et al.
@@
 In this work, we introduce a dynamic compensation method that continuously adjusts the pump frequency to match the instantaneous PIFS during the adiabatic initialization.
-We experimentally demonstrate the advantage of this dynamic compensation for a Kerr-cat qubit using a novel Nonlinearity-Engineered Triple-loop SQUID (NETS) device Hua et al.
-.
+We experimentally demonstrate the advantage of this dynamic compensation for a Kerr-cat qubit using a novel Nonlinearity-Engineered Multi-loop SQUID (NEMS) device Hua et al.
+(2024).
 This device not only holds promise for realizing bias-preserving CNOT gates Puri et al.
@@
 (2017, 2018); Sivak et al.
+(2019); He et al.
 (2023), ATS Lescanne et al.
@@
 (2012), and a series of asymmetric DC SQUIDs Iyama et al.
+(2024); Hoshi et al.
 (2024).
@@
 Figure 2: (a) Circuit schematic of the device.
-The Nonlinearity-Engineered Triple-loop SQUID (NETS) is shunted with a large capacitor, similar to that of a transmon.
-(b) Circuit schematic of the NETS.
+The Nonlinearity-Engineered Multi-loop SQUID (NEMS) is shunted with a large capacitor, similar to that of a transmon.
+(b) Circuit schematic of the NEMS.
 $I_{2}$ is the flux pumping port.
@@
 SM .
-(c) Photograph of the NETS.
+(c) Photograph of the NEMS.
 (d) Illustration comparing the dynamic and static compensation methods.
@@
 1(c), enabling high-fidelity initialization of the Kerr-cat qubit.
-Experiments.- The Kerr-cat qubit is implemented using a novel NETS device with its schematic circuit shown in Fig.
+Experiments.- The Kerr-cat qubit is implemented using a novel NEMS device with its schematic circuit shown in Fig.
 2(a).
 An ancillary transmon qubit is introduced for characterizing the Kerr-cat qubit.
-The NETS consists of three loops [Fig.
+The NEMS consists of three loops [Fig.
 2(b)], holding the advantage of independently tuning the nonlinearities of different orders by controlling the external flux through three current lines $I_{1}$, $I_{2}$, and $I_{3}$ Hua et al.
-; SM .
-Figure 2(c) depicts the fabricated NETS device, with the ratios of the junctions and loop areas being meticulously designed.
-By biasing the NETS at $(\phi_{\mathrm{e1}},\phi_{\mathrm{e2}},\phi_{\mathrm{e3}})=(-0.08,1.08,0)\pi$ and applying a flux pump via $I_{2}$ line, the Hamiltonian of Eq.
+(2024); SM .
+Figure 2(c) depicts the fabricated NEMS device, with the ratios of the junctions and loop areas being meticulously designed.
+By biasing the NEMS at $(\phi_{\mathrm{e1}},\phi_{\mathrm{e2}},\phi_{\mathrm{e3}})=(-0.08,1.08,0)\pi$ and applying a flux pump via $I_{2}$ line, the Hamiltonian of Eq.
 1 for the Kerr-cat qubit can be realized SM .
-Without an external pump, the NETS functions as a nonlinear oscillator at a frequency of $\omega_{\mathrm{a}}/2\pi=5.6\,\mathrm{GHz}$ with a Kerr nonlinearity $K/2\pi=6.9\,\mathrm{MHz}$.
+Without an external pump, the NEMS functions as a nonlinear oscillator at a frequency of $\omega_{\mathrm{a}}/2\pi=5.6\,\mathrm{GHz}$ with a Kerr nonlinearity $K/2\pi=6.9\,\mathrm{MHz}$.
 The amplitude damping and transverse coherence times are $T_{\mathrm{1N}}=6.0\,\mathrm{\mu s}$ and $T_{\mathrm{2N}}=3.0\,\mathrm{\mu s}$, respectively.
-The Kerr-cat qubit is realized by applying a flux pump at $\omega_{\mathrm{s}}/2\pi\approx 11.2\,\mathrm{GHz}$ with an amplitude of $\varepsilon_{\mathrm{2,KC}}/2\pi=9.7\,\mathrm{MHz}$ to the NETS.
-By initializing the NETS in the state $(\left|0\right\rangle+\left|1\right\rangle)/\sqrt{2}$, the Kerr-cat qubit can be prepared in the $\left|\alpha\right\rangle$ state on the equator of the Bloch sphere.
-The precession due to PIFS is measured by varying the holding time of NETS in the Kerr-cat states, allowing the calibration of $\Delta_{\mathrm{as,KC}}$ SM .
-Furthermore, the Rabi drive of the Kerr-cat qubit is realized by a charge drive on the NETS, with a Rabi frequency along the X-axis given by $\Omega_{X}=\mathrm{Re}(4\epsilon_{\mathrm{x}}\alpha)/\sqrt{1-\exp(-4|\alpha|^{2})}$  Puri et al.
+The Kerr-cat qubit is realized by applying a flux pump at $\omega_{\mathrm{s}}/2\pi\approx 11.2\,\mathrm{GHz}$ with an amplitude of $\varepsilon_{\mathrm{2,KC}}/2\pi=9.7\,\mathrm{MHz}$ to the NEMS.
+By initializing the NEMS in the state $(\left|0\right\rangle+\left|1\right\rangle)/\sqrt{2}$, the Kerr-cat qubit can be prepared in the $\left|\alpha\right\rangle$ state on the equator of the Bloch sphere.
+The precession due to PIFS is measured by varying the holding time of NEMS in the Kerr-cat states, allowing the calibration of $\Delta_{\mathrm{as,KC}}$ SM .
+Furthermore, the Rabi drive of the Kerr-cat qubit is realized by a charge drive on the NEMS, with a Rabi frequency along the X-axis given by $\Omega_{X}=\mathrm{Re}(4\epsilon_{\mathrm{x}}\alpha)/\sqrt{1-\exp(-4|\alpha|^{2})}$  Puri et al.
 (2020).
@@
 This can be explained by the collapse of the energy levels of the Fock states with respect to the compensated detunings, and the condition for the adiabatic evolution is violated.
-This leads to the observation of Landau-Zener transitions Landau (1932); Zener and Fowler (1932): an initial Fock state can be mapped to two stabilized states of the pumped NETS when the energy gap between the Fock state and other higher levels becomes small.
+This leads to the observation of Landau-Zener transitions Landau (1932); Zener and Fowler (1932): an initial Fock state can be mapped to two stabilized states of the pumped NEMS when the energy gap between the Fock state and other higher levels becomes small.
 Due to the energy difference of the stabilized states, the accumulated phase leads to an interference between the two states when mapping the states back to the Fock state basis, exhibiting the sharp oscillations.
@@
 The static compensation method shows comparable performance to the dynamic compensation approach when $K<\Delta_{\mathrm{as,KC}}<3K$.
-This can be attributed to the fact that a precisely symmetric ramping-down process can adiabatically map the stabilized NETS state back to the original Fock space, regardless of whether the NETS state is inside or outside the Kerr-cat qubit space.
+This can be attributed to the fact that a precisely symmetric ramping-down process can adiabatically map the stabilized NEMS state back to the original Fock space, regardless of whether the NEMS state is inside or outside the Kerr-cat qubit space.
 To directly characterize the initialization of Kerr-cat qubits, we perform a direct cat readout Grimm et al.
 (2020); Ding et al.
-(2024); SM of the NETS by coherently converting its state to the readout cavity.
+(2024); SM of the NEMS by coherently converting its state to the readout cavity.
 Measurements along different axes are realized with an X/2 gate implemented through Rabi drives and a Z/2 gate achieved by instantaneously switching off the flux pump Grimm et al.
@@
 (2024).
-We find that the fidelity degrades with $\Delta_{\mathrm{as,KC}}$ due to the increasing decoherence rates with the Kerr-cat size, and better performance is possible by improving the coherence of the NETS.
-For instance, a Kerr-cat qubit initialization with a fidelity exceeding 99% is anticipated when the NETS has $T_{\mathrm{1N}}=60\,\mathrm{\mu}$s and $T_{\mathrm{2N}}=30\,\mathrm{\mu}$s.
-Conclusion.- We have demonstrated high-fidelity initialization of a Kerr-cat qubit using a novel NETS device with a dynamic compensation method.
+We find that the fidelity degrades with $\Delta_{\mathrm{as,KC}}$ due to the increasing decoherence rates with the Kerr-cat size, and better performance is possible by improving the coherence of the NEMS.
+For instance, a Kerr-cat qubit initialization with a fidelity exceeding 99% is anticipated when the NEMS has $T_{\mathrm{1N}}=60\,\mathrm{\mu}$s and $T_{\mathrm{2N}}=30\,\mathrm{\mu}$s.
+Conclusion.- We have demonstrated high-fidelity initialization of a Kerr-cat qubit using a novel NEMS device with a dynamic compensation method.
 By carefully analyzing eigenstate evolution of the system under adiabatic pump ramping, we clarify the imperfections arising from PIFS and Landau-Zener transitions.
@@
 At a PIFS to Kerr nonlinearity ratio of $1.4$, the measured initialization fidelity improves from 57% to 78%, indicating a projected fidelity of 91% after excluding SPAM errors.
-Our demonstration with NETS and dynamic compensation paves the way for future advancements in scalable Kerr-cat qubits, enabling their application in quantum information processing and exploration of noise-biased fault-tolerant techniques.
+Our demonstration with NEMS and dynamic compensation paves the way for future advancements in scalable Kerr-cat qubits, enabling their application in quantum information processing and exploration of noise-biased fault-tolerant techniques.
 Acknowledgements.
@@
 11925404, 92165209, 92365301, 12061131011, 92265210, 11890704, 92365206), Innovation Program for Quantum Science and Technology (Grant No.
-2021ZD0300203 and 2021ZD0301800), and the National Key R&D Program (2017YFA0304303).
+2021ZD0300203, 2021ZD0301800, and 2021ZD0300201), and the National Key R&D Program (2017YFA0304303).
 This work was also supported by the Fundamental Research Funds for the Central Universities and USTC Research Funds of the Double First-Class Initiative.
 This work was partially carried out at the USTC Center for Micro and Nanoscale Research and Fabrication.
+Note added.- When finalizing this work, we became aware of a related study Iyama et al.
+(2024) that also utilized a chirped pulse for initialization.
+While that study focuses on the characterization and dynamics of cat states, our work delves into the physical mechanism of the adiabatic evolution of the state mapping process.